# SMART: Submodular Data Mixture Strategy for Instruction Tuning

### v2

@@ 4.9 Discussion
In Section 4.7, we saw that Graph Cut proves to be the most effective for selecting the weighted task subsets. Figures 678 of Appendix D contain t-SNE visualizations for task subsets selected by the three submodular functions in Table 1. Facility Location being a sum-max formulation, a single point is sufficient to represent a cluster and hence it gives more weightage to cluster centers and very less weightage to others. Log Determinant on the other hand predominantly selects diverse points not taking representation into account. Graph Cut, which models both, hence performs better than Facility Location and Log Determinant.
Further, in Figure 2 of Section 4.6, we saw that, at around 16 tasks, there is a peak in performance, after which there is a slight decline. To facilitate more investigation into what tasks are assigned more weightage, Table 5 and Table 6 list down the top-128 and last-128 tasks in the submodular ordering obtained using graph cut. While the former set contains more traditional NLP tasks like Natural Language Inference, Next Sentence Prediction, Question Answering, Summarization, etc ., the latter set mostly consists of tasks like Program Execution.
@@ 5.1 Data for Instruction Tuning
Research diverges on scaling instruction data quantity, with some advocating for limited data(Zhou et al., 2023; Chen et al., 2023a) to expose pretraining knowledge, while others argue for scaling up(Wei et al., 2021; Sanh et al., 2021). According to Ji et al. (2023); Dong et al. (2023); Yuan et al. (2023); Song et al. (2023), the impact of scaling varies across tasks and model abilities.
@@
High-quality data is crucial in instruction tuning(Chia et al., 2023; Ding et al., 2023; Zhou et al., 2023). Wang et al. (2023a) use perplexity to select suitable instructions generated by models, while Li et al. (2023a) employ the language model itself to augment and curate high-quality training examples to improve its performance. Cao et al. (2023) propose InstructionMining, utilizing natural language indicators to predict inference loss as a proxy for data quality without human intervention and select the best subset based on this. Chen et al. (2023b) introduce AlpaGasus, which uses a strong LLM to select high-quality subsets. Attendu and Corbeil (2023) employs dynamic data subset selection by filtering out unimportant samples during finetuning, based on an extended EL2N metric(Paul et al., 2021; Fayyaz et al., 2022). Taori et al. (2023) propose #InsTag to assess instruction diversity in SFT datasets using ChatGPT. Additionally, Wan et al. (2023) propose Explore-Instruct, utilizing LLMs to actively explore domain-specific spaces and gather diverse instruction-tuning data. Wu et al. (2023) select new data points that are distinct from existing ones in the model embedding space, augmenting the training dataset iteratively to enhance diversity within subsets.
@@
Many previous works show the benefit of scaling the number of tasks (Wei et al., 2021; Chung et al., 2022; Wang et al., 2022; Sanh et al., 2021). However, works like Iyer et al. (2022); Longpre et al. (2023) have also acknowledged that task balancing is also very important for effective instruction tuning. Dong et al. (2023) explore data composition across GSM8k, Code Alpaca, and ShareGPT datasets, finding differential impacts of data scaling on performance across abilities and propose a Dual-stage mixed fine-tuning strategy as a promising solution to activate multiple abilities efficiently. Ivison et al. (2022) identifies relevant multitask subsets based on the similarity between the pre-trained model’s representations, using a small amount of target task data. Yin et al. (2023) uses instruction representations for task selection, acting as a replay strategy to mitigate catastrophic forgetting and improve generalization in continual learning. Lou et al. (2023); Zhang et al. (2023) provide a survey of instruction tuning in general and Wang et al. (2023b) provide a detailed survey of data management for instruction tuning. Yue et al. (2023) construct math generalist models via instruction tuning on hybrid of chain-of-thought and program-of-thought rationales in math.
@@ 5.2 Submodularity for Subset Selection
Submodularity (Fujishige, 2005) has a long history in combinatorial optimization, game theory, economics etc (Edmonds, 1970; Lovász, 1983; Carter, 2001; Topkis, 1998). It has recently gained traction in machine learning where it has been used for data subset selection for machine translation (Kirchhoff and Bilmes, 2014), speech recognition (Wei et al., 2014; Mittal et al., 2022; Kothawade et al., 2023), pre-training of LLMs (Renduchintala et al., 2023), active learning (Wei et al., 2015; Kothawade et al., 2021), hyperparameter tuning (Killamsetty et al., 2022), domain adaptation (Karanam et al., 2022), computer vision (Kaushal et al., 2019), continual learning (Tiwari et al., 2022) etc. For a more detailed review of submodularity and its applications, please refer to the survey by Bilmes (2022).
@@ 6 Conclusion & Future Work
This paper introduces a novel data mixture strategy for instruction tuning that utilizes a submodular function to assign importance scores to tasks, determine the mixture weights, and also select non-redundant samples from each task. Further, we reveal that in a low-budget setting, splitting the budget among a small subset of representative tasks yields superior performance when compared to dividing it among all tasks, which suggests that task scaling should be done more judiciously. Future work could explore making this method more model-specific (e.g., modify task embedding computation), modify approach for targeted instruction-data selection for creating expert LLMs.
@@ Appendix A Code and Data
Our code for creating data mixtures is open-sourced at https://anonymous.4open.science/r/SMART-D829. The code development utilized open-source tools, primarily relying on the HuggingFace library for model training with PyTorch as the underlying framework. Both PyTorch and HuggingFace are licensed under permissive licenses, with PyTorch under the BSD license and HuggingFace under the Apache 2.0 license. Additionally, submodular optimization was performed using SUBMODLIB, which is an openly accessible library on GitHub at https://github.com/decile-team/submodlib under the MIT license.
@@ Appendix B Task Scaling Curves: Varying $f_{1}$ and $f_{2}$
Figure 3: Task Scaling Curves for various $f_{1}$ and $f_{2}$ combinations
@@ Appendix C Task Scaling Curves: Mistral, Falcon
$N^{\prime}$ Data Mix. Mistral Falcon
$25000$ EPM(Baseline-1) 52.69 23.09
EM(Baseline-2) 51.65 21.1
SMART 53.33 22.73
$50000$ EPM(Baseline-1) 53.75 19.86
EM(Baseline-2) 53.16 22.96
SMART 53.8 26.31
$100000$ EPM(Baseline-1) 54.24 20.88
EM(Baseline-2) 54.57 24.85
SMART 53.75 27.22
$200000$ EPM(Baseline-1) 55.25 24.7
EM(Baseline-2) 54.71 24.39
SMART 55.4 30.96
Table 4: Comparison of SMART with baselines for Mistral and Falcon.

### v3

@@ 4.10 Discussion
In Section 4.8, we saw that Graph Cut proves to be the most effective for selecting the weighted task subsets. Figures 678 of Appendix D contain t-SNE visualizations for task subsets selected by the three submodular functions listed in Table 1. Facility Location being a sum-max formulation, a single point is sufficient to represent a cluster and hence it gives more weightage to cluster centers and very less weightage to others. As a result, the submodular gains for first few selected points are very high and then the gains quickly become very small for other points in case of facility location. Log Determinant on the other hand predominantly selects diverse points not taking representation into account. Graph Cut, which models both, hence performs better than both Facility Location and Log Determinant for selecting tasks.
Further, in Figure 1 of Section 4.6 we saw that, at around 16 tasks, there is a peak in performance, after which there is a slight decline. To facilitate more investigation into what tasks are assigned more weightage, Table 6 and Table 7 list down the top-128 and last-128 tasks in the submodular ordering obtained using graph cut. While the former set contains more traditional NLP tasks like Natural Language Inference, Next Sentence Prediction, Question Answering, Summarization, etc ., the latter set mostly consists of tasks like Program Execution.
@@ 5.1 Data for Instruction Tuning
Research diverges on scaling instruction data quantity, with some advocating for limited data (Zhou et al., 2023; Chen et al., 2023a) to expose pretraining knowledge, while others argue for scaling up (Wei et al., 2021; Sanh et al., 2021). According to Ji et al. (2023); Dong et al. (2023); Yuan et al. (2023); Song et al. (2023), the impact of scaling varies across tasks and model abilities. AlShikh et al. (2023) also introduce a metric for instruction following ability, and suggest an early stopping criterion for instruction-tuning.
@@
High-quality data is crucial in instruction tuning (Chia et al., 2023; Ding et al., 2023; Zhou et al., 2023). Wang et al. (2023a) use perplexity to select suitable instructions generated by models, while Li et al. (2023a) employ the language model itself to augment and curate high-quality training examples to improve its own performance. Cao et al. (2023) propose InstructionMining, utilizing natural language indicators to predict inference loss as a proxy for data quality without human intervention and select the best subset based on this. Chen et al. (2023b) introduce AlpaGasus, which uses a strong LLM to select high-quality subsets. Lu et al. (2023) and Madaan et al. (2024) also leverage the power of fine-tuned LLM itself to evaluate the quality of instructions. Attendu and Corbeil (2023) employs dynamic data subset selection by filtering out unimportant samples during finetuning, based on an extended EL2N metric (Paul et al., 2021; Fayyaz et al., 2022). Taori et al. (2023) propose #InsTag to assess instruction diversity in SFT datasets using ChatGPT. Additionally, Wan et al. (2023) propose Explore-Instruct, utilizing LLMs to actively explore domain-specific spaces and gather diverse instruction-tuning data. Wu et al. (2023) select new data points that are distinct from existing ones in the model embedding space, augmenting the training dataset iteratively to enhance diversity within subsets.
@@
Many previous works show the benefit of scaling the number of tasks (Wei et al., 2021; Chung et al., 2022; Wang et al., 2022; Sanh et al., 2021). However, works like Iyer et al. (2022); Longpre et al. (2023) have also acknowledged that task balancing is also very important for effective instruction tuning. Dong et al. (2023) explore data composition across GSM8k, Code Alpaca, and ShareGPT datasets, finding differential impacts of data scaling on performance across abilities and propose a Dual-stage mixed fine-tuning strategy as a promising solution to activate multiple abilities efficiently. Ivison et al. (2022) identifies relevant multitask subsets based on the similarity between the pre-trained model’s representations, using a small amount of target task data. Yin et al. (2023) uses instruction representations for task selection, acting as a replay strategy to mitigate catastrophic forgetting and improve generalization in continual learning. Yue et al. (2023) construct math generalist models via instruction tuning on hybrid of chain-of-thought and program-of-thought rationales in math. Lou et al. (2023); Zhang et al. (2023) provide a survey of instruction tuning in general and Wang et al. (2023b) provide a detailed survey of data management for instruction tuning.
@@ 5.2 Submodularity for Subset Selection
Submodularity (Fujishige, 2005) has a long history in combinatorial optimization, game theory, economics etc (Edmonds, 1970; Lovász, 1983; Carter, 2001; Topkis, 1998). It has recently gained traction in machine learning where it has been used for data subset selection for machine translation (Kirchhoff and Bilmes, 2014), speech recognition (Wei et al., 2014; Mittal et al., 2022; Kothawade et al., 2023), efficient pre-training of language models (Renduchintala et al., 2023), active learning (Wei et al., 2015; Kothawade et al., 2021), hyperparameter tuning (Killamsetty et al., 2022), domain adaptation (Karanam et al., 2022), computer vision (Kaushal et al., 2019), continual learning (Tiwari et al., 2022) etc. For a more detailed review of submodularity and its applications, please refer to the survey by Bilmes (2022).
@@ 6 Conclusion & Future Work
In this paper, we introduced SMART — a novel data mixture strategy for instruction tuning that utilizes a submodular function to assign importance scores to tasks, determine the mixture weights, and also select non-redundant samples from each task. Further, we also reveal that in a low-budget setting, splitting the budget among a small subset of representative tasks yields superior performance when compared to dividing it among all tasks, which suggests that task scaling should be done more judiciously. Future work could explore making this method more model-specific (e.g., modify task embedding computation) and also possibly modify the approach for targeted instruction-data selection for creating expert LLMs that specialize in specific skills such as math, code etc.
@@ Appendix A Code and Data
Our code we used for instruction tuning and for creating data mixtures is open-sourced at https://github.com/kowndinya-renduchintala/SMART. The code development utilized open-source tools, primarily relying on the HuggingFace library for model training with PyTorch as the underlying framework. Both PyTorch and HuggingFace are licensed under permissive licenses, with PyTorch under the BSD license and HuggingFace under the Apache 2.0 license. Additionally, submodular optimization was performed using SUBMODLIB, which is an openly accessible library on GitHub at https://github.com/decile-team/submodlib under the MIT license.
@@ Appendix B Task Scaling Curves: Varying $f_{1}$ and $f_{2}$
Figure 3: Task Scaling Curves for various $f_{1}$ and $f_{2}$ combinations for Llama-2-7b